OpenStudy (anonymous):

A bumblebee buzzes past you at 3 m/s. The frequency of the hum made by its wings is 152 Hz. Assume the speed of sound to be 342 m/s. What is the observed frequency of the hum as the bumblebee approaches you? 155 Hz 153 Hz 150 Hz 149 Hz

OpenStudy (anonymous):

\[\lambda = \frac{ v }{ f }\] \[\lambda = 342/152 = 2.25 m\] now you substract the bee speed from sound speed, because they are moving in the same direction 342-3=339m/s \[f=\frac{ v }{ \lambda } = \frac{ 339 }{ 2,25 } = 150,67 \] So the answer is 150,7 Hz